# Sustainability and Carbon Emissions of Future Accelerators

### v2

@@ 1 INTRODUCTION
However, over these same decades, our planet will be threatened by human-induced climate change at a scale that could disrupt our scientific plans. Should global warming exceed 1.5C there will be significant threats to human well-being due to reduced food and water supplies, increased rates of disease, and decreased availability of land due to sea level rise [3]. The societal challenges arising from this so-called “climate emergency” [4] could draw resources away from fundamental research, no matter how compelling the scientific case. The sixth assessment report of the Intergovernmental Panel on Climate Change has concluded that it is “unequivocal” that global warming is due to the emission of carbon dioxide and other greenhouse gases [5], and that to prevent warming beyond 1.5C, we must limit the emissions of these gases in the future. In particular, to achieve an 83% probability that warming will be limited to less than 1.5C, the total carbon budget for the planet is 300 gigatonnes of CO2e [^1] Carbon dioxide equivalent or CO2e means the number of metric tons of CO2 emissions with the same global warming potential as one metric ton of another greenhouse gas, and is calculated using Equation A-1 in 40 CFR (U.S. Code of Federal Regulations Part 98.). We use the metric ton (t = tonne) throughout this article., or about 1.1 t CO2e per year per capita until the year 2050, after which we must reach net zero carbon emissions worldwide. Current per capita emissions are 4.7 t CO2e worldwide (and also in France and the UK specifically), about 8 t CO2e in Germany, China and Japan, and 14.9 t CO2e in the United States [6]. Significant societal action must be taken, starting now, to meet this threat to our current way of life.
@@
In this article, we review the current understanding of the sustainability of future accelerator projects  [^2] A discussion of emissions associated with current particle physics projects and associated activities, as well as various HEP calculators, can be found in [7, 8, 9]., and highlight steps that can be taken to reduce their associated carbon emissions. We devote much of our consideration to the accelerators themselves, namely future energy-frontier colliders, in Section 2, as we expect that their construction and operations will lead to the largest amount of CO2e emissions in the life cycle of an accelerator project. We also provide briefer summaries of the emissions that might be associated with the operation of potential future particle detectors (Section 3), the computing facilities required to process and analyze the collected experimental data (Section 4), and travel and other activities necessary to carry out research in large international collaborations (Section 5). We conclude with suggestions on potential lines of research on mitigating greenhouse gas emissions in particle physics (Section 6) and closing thoughts (Section 7).
@@ 2 EMISSIONS ASSOCIATED WITH FUTURE ENERGY-FRONTIER COLLIDERS
We consider the carbon emissions associated with ten possible future energy-frontier colliders, which all address physics goals identified in the most recent P5 [1] and European Strategy [13] reports. Some of the colliders are explicitly mentioned in those reports, while others have not received as much attention. We make assumptions on their location based on siting plans described by the individual projects. These are:
@@
  • Future Circular Collider (FCC) [18]: The FCC would be the largest accelerator ever built. A 90.2 km circumference tunnel would be constructed, sited to connect to the current CERN complex. The initial phase, FCC-ee [19], would provide $e^{+}e^{-}$ collisions from 88 to 365 GeV CM energy. After that 15-year physics program is completed, a hadron collider, FCC-hh [20], would be built in the same tunnel, with the goal of reaching 100 TeV CM proton collisions, which would yield approximately 10 TeV parton center-of-momentum (pCM) collisions.
@@ 2.1 Civil Construction
Table 1 gives our tabulation of emissions associated with each of the tunnel and cavern construction for the ten collider projects considered. As described below, several of the future collider projects are now collaborating with external firms to provide estimates of emissions, or undertaken their own calculations, using the LCA framework. Wherever possible, we make use of the projects’ own estimates, and for projects that have not completed their own assessments, we have leveraged results from other projects to create similar estimates.
@@
C3 has completed an estimate of their emissions from construction, following the ARUP procedure [33]. Their analysis yields an estimate of the emissions per length of tunnel of 17 kt CO2e/km, about a factor of two greater than for CLIC, suggesting that further optimizations can be completed, but for calculating the total emissions this is compensated for by the much shorter tunnel length.
@@ 2.2 Accelerator Components
All of the above estimates only consider the global warming potential associated with the construction of tunnels and other auxiliary facilities, and most importantly the production of steel and concrete associated with them. We also consider the carbon impacts due to the procurement of the accelerator components themselves, the many magnets and RF cavities that are necessary for an energy-frontier collider. In a linear collider the dominant structures are the superconducting RF cavities. As an example and a crude estimate, we consider the cavities needed for the ILC, which are primarily made of niobium. Niobium has very good conductivity, making it suitable for a SRF cavity, but it is designated a “critical” element on e.g. the U.S. draft list of critical minerals, which ranked niobium second out of 50 critical mineral commodities [34]. A high criticality rating indicates that a “mineral commodity has important strategic (e.g., national security) end uses, but its supply is insecure due to various socioeconomic, geological, and geopolitical factors causing increased supply risk”. Niobium is increasingly used in transport, energy generation and storage and as such demand for niobium is expected to grow in line with countries’ targets toward net zero carbon emissions. Based on a description of the cavity geometry [15] we estimate that each cavity includes a mass of 39.6 kg of niobium, which has a carbon intensity of about 75 kg CO2e/kg  [^3] This carbon intensity number is obtained from ChatGPT, which quotes a value between 50 and 100 kg CO2e/kg based on general knowledge of the energy-intensive nature of niobium extraction and refinement processes. We could not find a LCA for pure niobium with high Residual Resistivity Ratio used for SRF cavities; most LCA analyses are for ferroniobium and niobium oxides [35].. With 8000 cavities expected in the 250 GeV ILC, the total emissions associated with the niobium is about 26 kt CO2e, which is about 27% of the contribution from the civil construction. The cavities therefore have subleading but still significant emissions.
For a circular collider, such as the FCC-ee, the dominant elements of the accelerator are the magnets. The 2900 dipole magnets required have 219 kg/m of iron and 19.9 kg/m of aluminum and each dipole has a magnetic length of 24 m [36]. Assuming a carbon intensity of 2.0 kg CO2e/kg and 6.8 kg CO2e/kg for iron and aluminum, respectively [27], the total emissions associated with the materials is about 40 kt CO2e. Although this is large in absolute terms, it is about 7% of the emissions associated with the civil construction, hence also a subleading factor.
@@ 2.3 Operations
Power consumption has long been a concern for accelerator design, as it has historically been the most financially costly element of accelerator operations. As we aim to operate a sustainable accelerator, we must also consider the cost to the climate due to the carbon emissions associated with power consumption. In doing so, we consider the broader context of how power will be generated on the timescale of the operation of future colliders. All of the projects described in the previous section are at least ten years away from the start of operations, if not significantly more, given the realities of the world’s existing collider programs that are not yet completed and the availability of funding for new construction [1]. The Paris Agreement [37] mandates a 43% decline in greenhouse gas emissions by 2030, far earlier than the operation of any future colliders, which can only be achieved through a substantial de-carbonization of electrical power production.
@@
Table 2: Carbon emissions due to operations of future colliders using the carbon intensity from the APS scenario of the IEA. The total electrical ${}_{\mbox{e}}$) corresponding to the duration of operations for each respective collider is also shown. For the CCC the numbers in brackets correspond to the optimized power design [33]. A * is inserted to highlight that this value is in the context of a scope analysis [10].
@@
Figure 1: Emissions of future energy-frontier colliders split according to civil construction, accelerator components and operations. The SRF cavitites emissions correspond to an estimate for the ILC while the dipole magnets come from the FCC-ee design. Colliders located in Europe or the USA have zero emissions operations from 2040, following the Scope 2 convention defined in [10] and the APS scenario from the IEA. See text for more information on all numbers.
@@ 3 EMISSIONS ASSOCIATED WITH POTENTIAL FUTURE DETECTORS
Energy-frontier detectors that will operate after 2040 will very likely not have access to HFCs or other fluorinated gases, as there is an ongoing phase-down of such gases [46]. There is already very active R&D in particle physics for finding so-called eco-gas replacements to the GHG used in detectors. For example, at the Gamma Irradiation Facility (GIF) the RPC ECOGas@GIF++ collaboration is performing ageing tests using muon beams [47]. In general, hydrofluoroolefin (HFO) and perfluoroketone (PKF) are promising molecules to replace GHG. For example, the gas made out of C3H2F4/R-1234ze is a promising replacement to the C2H2F4/R-134a used in RPCs. This gas has a GWP of less than seven. Studies done so far have found that in general a higher high-voltage working point of the detector is needed and higher currents are observed [48]. Unfortunately, some studies have also shown some significant ageing issues for detectors using such a gas as part of their mixture [49].
@@ 4 EMISSIONS ASSOCIATED WITH COMPUTING FOR FUTURE PROJECTS
Given the expectation of nearly carbon-free power generation on the timescale of future colliders, we anticipate that the carbon impact of operating the computing resources for these experiments will be negligible in comparison to that of the construction of the collider and detector infrastructure. Instead, other concerns will come into play. For instance, as discussed in Section 6, as human activity becomes increasingly electrified as a means to reduce the use of fossil fuels, the supply of electricity may be increasingly constrained compared to the overall demands. Therefore, a focus on energy efficiency of computing hardware, and the software algorithms that run on them, will still be needed, even if the energy used is carbon-free. This can be achieved in part through increasing parallelization of processing and a transition of processing hardware from traditional central processing units (CPUs) to graphical processing units (GPUs) and tensor processing units (TPUs). Studies of science application performance at modern supercomputing facilities have demonstrated significant improvements in energy efficiency by using GPUs instead of CPUs for processing, although the impacts vary significantly with the specific application [58]. However, history has repeatedly demonstrated that increased efficiency does not necessarily translates into reduced total energy (Jevons paradox) [59], including in the case of the usage of computing farms.
The dominant concern in the era of de-carbonized power generation will be the global warming impacts associated with the manufacturing process that leads to “embodied carbon” in the computing hardware. This involves the emissions associated with extracting the raw materials (such as heavy metals) that are needed in computers, and the emissions associated with the manufacturing process itself. For instance, the embodied carbon associated with a typical Dell server used in a data center is about 1300 kgCO2[60]. The embodied carbon associated with computer hardware can be reduced through improvements to manufacturing processes, the use of lower-impact materials, and an improved recycling practice of computing components [61]. While these are beyond the control of the end users, particle physicists can seek to reduce the impact of embodied carbon by selecting vendors with the most sustainable practices, by trying to extend the lifetime of existing computing hardware, thereby reducing the rate of replacement and by adhering to strict recycling practices.
@@ 5 EMISSIONS ASSOCIATED WITH TRAVEL AND OTHERS
We use this information as inputs to a simple model (see Table 4) of the carbon emissions associated with travel to the experiment site by collaborators, using the FCC as an example. We assume a collaboration size of 5000, in which half of the collaborators reside near CERN, the experiment site. One quarter are “regional” participants who work London-like distances from CERN and visit the experiment four times per year. The remaining quarter are “remote” participants who live Chicago-like distances from CERN and visit the experiment only two times per year, because of the large distance. The regional participants produce about 1.7 kg CO2e/year from their air travel, while the remote participants produce about 8.4 kg CO2e/year. In one year, the collaboration produces 12.6 kt CO2e emissions from air travel, which summed over a 20-year physics program is 0.25 Mt CO2e. Comparing to the estimates of Section 2, this amount is more than that associated with the FCC accelerator components, and in fact about one quarter of the emissions associated with the civil construction, a subleading but still significant contribution to the carbon emissions over the lifetime of the project.
@@
International conferences also has the potential to be a significant source of carbon emissions, as they draw people from all over the planet, many of whom arrive by long-haul airplane flights. This is issue is not unique to future colliders, and we refer readers to previous discussions of the topic [67].
@@ 6 OBSERVATIONS AND R&D PRIORITIES
Research and development activities focused on minimizing the environmental impact of future colliders have been in progress for a few years. Several presentations in the 2024 editions of the Sustainable HEP conference [68] and of the ICHEP conference [69] reported on those efforts. Of note was the announcement of the start of the iSAS (innovate for Sustainable Accelerating Systems) [70] project funded by EU Horizon 2020. This initiative plans to explore three key innovation directions for energy savings: in the RF system, the cryogenics, and energy savings from the beam itself [71]. In addition, the CEPC collaboration showed interesting results regarding the power savings which could be achieved using the mid-temperature baking technology, which can enhance the quality factor of the RF cavities by a factor of five compared to the current electropolishing technique [30]. Finally, more projects are using the LCA method to optimize their designs in view of reducing their environmental impact, as demonstrated by the ISIS-II project [72].
The accelerator construction analysis described in Section 2.1 suggests that the emissions associated with the cement needed for the tunnel could easily dominate the total emissions of a future accelerator, and thus is in the greatest need of mitigation. Consequently, any new accelerator design which minimizes the length of the tunnels has a clear advantage from a carbon emissions and potentially a cost perspective. The recent hybrid, asymmetric, linear Higgs factory design based on plasma-wakefield and RF acceleration is such an example [73]. As mentioned by the engineering companies involved in the various LCA estimates available, the choice of material used for the civil construction can result in a wide range of embedded carbon. As an example, in the standard ICE database [27] there are 23 different types of cement, which is responsible for the high embedded carbon of concrete. The standard cement assumed by most LCAs is CEM I, Ordinary Portland Cement, with a carbon intensity of 0.912 kgCO2e/kg. As explained in [29] if this cement is mixed in with some Supplementary Cementitious Materials (SCMs), such as fly ash, Ground Granulated Blast furnace Slag (GGBS), limestone powder, or calcined clay, then the carbon intensity may be reduced. For example, the CEM II/B-S - 28% GGBS which includes between 21-25% of GGBS material has a carbon intensity of 0.672 kgCO2e/kg, a reduction of 26% compared to CEM I. However, one has to consider the potential drawbacks of using those alternative materials including whether it would be appropriate for a structure that is meant to have a lifetime greater than sixty years and what the cost implications would be, given that those materials are becoming much more in demand due to sustainability concerns worldwide. Finally, there are also very important R&D avenues being investigated in sequestering carbon during the cement production stage [74].
@@
As discussed in section 4, emissions from computing originate from the embedded carbon in the computing components and from the electricity used to power those components. For the latter, as mentioned in section 2.3, western countries expect to have a decarbonized electricity grid by 2040, hence a significant fraction of computing farms that will be used to process the energy-frontier collected data will have negligible indirect emissions. However, as already mentioned, demand for electricity is expected to grow significantly due to the necessity for countries to achieve their Net Zero target by 2050. For example, in the U.K., the Climate Change Committee forecasts that electricity demand will double by 2050 compared to 2018 levels [32], with a similar picture in the U.S. [76]. In addition, installed capacity needs to be at least a factor of two larger than the peak demand for electricity. Achieving this target while only using renewable or nuclear sources will be a significant challenge and consequently it is reasonable to assume that the price of electricity will either remain the same as today or more likely increase. As a consequence, it is of vital importance for computing farms and for colliders used for particle physics projects to use as little power and total electrical energy as possible. An interesting area of computing R&D studies is frequency modulation where the CPU clock can run at different rates and be dynamically adjusted to adapt to the power demand of the grid [77].
@@
Some authors have suggested that one possible avenue to ensure that renewable electricity sources are used for the power needs of future facilities is for those facilities to purchase such renewable providers (e.g. solar farms or wind turbine farms). It has also been suggested that CERN could get its electricity from solar farms in North Africa, transmitted via sub-sea cables (see Case Study 3.3 in [7]). First it is important to note that Scope 2 (indirect) emissions do not include the emissions associated with the construction and the embodied carbon of the electricity sources, while if a facility builds or purchases such a source, those emissions would need to be included in their Scope 1 emissions. Finally, we believe that the implications of particle physics research facilities acquiring such resources should be explored and discussed. Given the climate emergency facing humanity, particle physics researchers should instead contribute to the work needed for every country to decarbonize their electricity grid, in turn ensuring the minimization of Scope 2 emissions for particle physics projects.
@@
As discussed in section 5, one way to mitigate travel emissions is to maximize the use of videoconferencing systems. However, as experienced during the COVID-19 pandemic, this does not capture the spontaneous interactions crucial to the exchange of ideas, and it can also lead to online fatigue. The field of particle physics has a long tradition of being an innovator in collaborative tools, having for example invented the World Wide Web. This spirit of innovation should continue in order to solve the remaining hurdles in the way of minimum travel. This will be especially crucial given that decarbonizing flight is a very difficult goal to achieve and some countries have Green Parties advocating legislation aimed at introducing a frequent-flyer levy [78]. The field of particle physics could for example champion a vast expansion of the use of Virtual/Augmented Reality headsets which can help make online social interactions more natural and also ensures participation from the Global South in particle physics projects.

### v3

@@ 1 INTRODUCTION
However, over these same decades, our planet will be threatened by human-induced climate change at a scale that could disrupt our scientific plans. Should global warming exceed 1.5C there will be significant threats to human well-being due to reduced food and water supplies, increased rates of disease, and decreased availability of land due to sea level rise [3]. The societal challenges arising from this so-called “climate emergency” [4] could draw resources away from fundamental research, no matter how compelling the scientific case. The sixth assessment report of the Intergovernmental Panel on Climate Change has concluded that it is “unequivocal” that global warming is due to the emission of carbon dioxide and other greenhouse gases [5], and that to prevent warming beyond 1.5C, we must limit the emissions of these gases in the future. In particular, to achieve an 83% probability that warming will be limited to less than 1.5C, the total carbon budget for the planet is 300 gigatonnes of CO2e [^1] Carbon dioxide equivalent or CO2e is the number of metric tons of CO2 emissions with the same global warming potential as one metric ton of another greenhouse gas. It is calculated using Equation A-1 in U.S. Code of Federal Regulations Title 40, Section 98.2 [6]. We use the metric ton (t = tonne) throughout this article., or about 1.1 t CO2e per year per capita until the year 2050, after which we must reach net zero carbon emissions worldwide. Current per capita emissions are 4.7 t CO2e worldwide (and also in France and the UK specifically), about 8 t CO2e in Germany, China and Japan, and 14.9 t CO2e in the United States [7]. Significant societal action must be taken, starting now, to meet this threat to our current way of life.
@@
In this article, we review the current understanding of the sustainability of future accelerator projects  [^2] A discussion of emissions associated with current particle physics projects and associated activities, as well as various HEP calculators, can be found in [8, 9, 10]., and highlight steps that can be taken to reduce their associated carbon emissions. We devote much of our consideration to the accelerators themselves, since we expect that their construction and operation will lead to the largest amount of CO2e emissions in the life cycle of an accelerator project. In Section 2, we will explain the factors that contribute to the CO2e emissions through explicit estimates for a number of proposed colliders. We also provide briefer summaries of the emissions that might be associated with the operation of potential future particle detectors (Section 3), the computing facilities required to process and analyze the collected experimental data (Section 4), and travel and other activities necessary to carry out research in large international collaborations (Section 5). We conclude with suggestions on potential lines of research on mitigating greenhouse gas emissions in particle physics (Section 6) and closing thoughts (Section 7).
While our focus is on particle accelerators used for high-energy physics research, we note that accelerators are used for many other applications in industry, medicine, chemical and biological research, and defense. Improvements in sustainability in particle physics might also provide benefits to other fields.
@@ 2 EMISSIONS ASSOCIATED WITH FUTURE ENERGY-FRONTIER COLLIDERS
We consider the carbon emissions associated with ten possible future energy-frontier colliders, which all address physics goals identified in the most recent P5 [1] and European Strategy [14] reports. Some of the colliders are explicitly mentioned in those reports, while others have not received as much attention. We make assumptions on their location based on siting plans described by the individual projects. We note that while we are following assumptions of the conditions during construction and operation as currently proposed, these projects are all still evolving. It is entirely possible that improvements in carbon intensity will be made, especially if the research and development activities described in Section 6 are pursued. Our evaluation is only a summary of the status at the time of writing. The projects are:
@@
  • Future Circular Collider (FCC) [19]: The FCC would be the largest accelerator ever built. A 90.2 km circumference tunnel would be constructed, sited to connect to the current CERN complex. The initial phase, FCC-ee [20], would provide $e^{+}e^{-}$ collisions from 88 to 365 GeV CM energy. After that 15-year physics program is completed, a hadron collider, FCC-hh [21], would be built in the same tunnel, with the goal of reaching 100 TeV CM proton collisions. This would yield approximately 10 TeV parton center-of-momentum (pCM) collisions.
@@ 2.1 Civil Construction
Table 1 gives our tabulation of emissions associated with each of the tunnel and cavern construction for the ten collider projects considered. As described below, several of the future collider projects are now collaborating with external firms to provide estimates of emissions, or undertaken their own calculations, using the LCA framework. Wherever possible, we make use of the projects’ own estimates, and for projects that have not completed their own assessments, we have leveraged results from other projects to create similar estimates. Ultimately, each project will need to complete their own detailed LCA assessments, led by expert engineers in collaboration with project physicists.
@@
C3 has completed an estimate of their emissions from construction, following the ARUP procedure [34]. Their analysis yields an estimate of the emissions per length of tunnel of 17 kt CO2e/km, about a factor of two greater than for CLIC, suggesting that further optimizations can be completed. For calculating the total emissions this is compensated for by the much shorter tunnel length.
@@ 2.2 Accelerator Components
All of the above estimates only consider the global warming potential associated with the construction of tunnels and other auxiliary facilities, and most importantly the production of steel and concrete associated with them. We also consider the carbon impacts due to the procurement of the accelerator components themselves, the many magnets and RF cavities that are necessary for an energy-frontier collider.
In a linear collider the dominant structures are the superconducting RF cavities. As an example and a crude estimate, we consider the cavities needed for the ILC, which are primarily made of niobium. Niobium has very good conductivity, making it suitable for a SRF cavity, but it is on the U.S. draft list of critical minerals, which ranked niobium second out of 50 critical mineral commodities [35]. A high criticality rating indicates that a “mineral commodity has important strategic (e.g., national security) end uses, but its supply is insecure due to various socioeconomic, geological, and geopolitical factors causing increased supply risk”. Niobium is increasingly used in transport, energy generation and storage and as such demand for niobium is expected to grow in line with countries’ targets toward net zero carbon emissions. Based on a description of the cavity geometry [16] we estimate that each cavity includes a mass of 39.6 kg of niobium.
The pure niobium needed for these cavities requires significant processing to achieve a low residual resistivity ratio [15, 36]. Information on the carbon intensity of mining and producing pure niobium is limited, although there have been LCA studies on ferroniobium and niobium oxides [37]. However, the mining and processing of niobium should be similar to that for titanium. The carbon intensity of titanium alloys is of the order of 50 kg CO2e/kg [38]. Although this is a very large carbon intensity compared to other materials mentioned, it is not as large as that of precious metals used in catalytic converters and electronic components, such as palladium (8,500 kg CO2e/kg) and platinum (15,000 kg CO2e/kg). For our calculations we estimate a carbon intensity of 75 kg CO2e/kg for niobium.
With 8000 cavities expected in the 250 GeV ILC, the total emissions associated with the niobium is about 26 kt CO2e, which is about 10% of the contribution from the total civil construction. The cavities therefore have subleading but still significant emissions.
For a circular collider, such as the FCC-ee, the dominant elements of the accelerator are the magnets. The 2900 dipole magnets required have 219 kg/m of iron and 19.9 kg/m of aluminum and each dipole has a magnetic length of 24 m [39]. Assuming a carbon intensity of 2.0 kg CO2e/kg and 6.8 kg CO2e/kg for iron and aluminum, respectively [28], the total emissions associated with the materials is about 40 kt CO2e. Although this is large in absolute terms, it is about 4% of the emissions associated with the total civil construction, hence also a subleading factor.
@@ 2.3 Operations
Power consumption has long been a concern for accelerator design, as it has historically been the most financially costly element of accelerator operations. As we aim to operate a sustainable accelerator, we must also consider the cost to the climate due to the carbon emissions associated with power consumption. In doing so, we consider the broader context of how power will be generated on the timescale of the operation of future colliders. All of the projects described in the previous section are at least ten years away from the start of operations. This may be an underestimate, given that the world’s existing collider programs are not yet completed and the limited availability of funding for new construction [1]. The Paris Agreement [40] mandates a 43% decline in greenhouse gas emissions by 2030, far earlier than the operation of any future colliders, which can only be achieved through a substantial de-carbonization of electrical power production.
@@
Table 2: Carbon emissions due to operations of future colliders using the carbon intensity from the APS scenario of the IEA. The total electrical ${}_{\mbox{e}}$) corresponding to the duration of operations for each respective collider is also shown. For the CCC the numbers in brackets correspond to the optimized power design [34]. A * is inserted to highlight that this value is in the context of a scope analysis [11]. In the last column it is assumed that electrical power is decarbonized in the U.S. and in Europe, as described in the text.
@@
Figure 1: Emissions of future energy-frontier colliders split according to civil construction, operations and accelerator components. The accelerator components included for representative purposes are the SRF cavitites for the ILC and the dipole magnets from the FCC-ee design. Colliders located in Europe or the U.S. have zero emissions operations from 2040, following the Scope 2 convention defined in [11] and the APS scenario from the IEA. See text for more information on all numbers.
@@ 3 EMISSIONS ASSOCIATED WITH POTENTIAL FUTURE DETECTORS
Energy-frontier detectors that will operate after 2040 will very likely not have access to HFCs or other fluorinated gases, as there is an ongoing phase-down of such gases [49]. There is already very active R&D in particle physics for finding so-called eco-gas replacements to the GHG used in detectors. For example, at the Gamma Irradiation Facility (GIF) the RPC ECOGas@GIF++ collaboration is performing aging tests using muon beams [50]. In general, hydrofluoroolefin (HFO) and perfluoroketone (PKF) are promising molecules to replace GHG. For example, the gas made out of C3H2F4/R-1234ze is a promising replacement to the C2H2F4/R-134a used in RPCs. This gas has a GWP of less than seven. Studies done so far have found that in general a higher high-voltage working point of the detector is needed and higher currents are observed [51]. Unfortunately, some studies have also shown some significant aging issues for detectors using such a gas as part of their mixture [52].
@@ 4 EMISSIONS ASSOCIATED WITH COMPUTING FOR FUTURE PROJECTS
Given the expectation of nearly carbon-free power generation on the timescale of future colliders, we anticipate that the carbon impact of operating the computing resources for these experiments will be negligible in comparison to that of the construction of the collider and detector infrastructure. Instead, other concerns will come into play. For instance, as discussed in Section 6, as human activity becomes increasingly electrified as a means to reduce the use of fossil fuels, the supply of electricity may be increasingly constrained compared to the overall demands. Therefore, a focus on energy efficiency of computing hardware, and the software algorithms that run on them, will still be needed, even if the energy used is carbon-free. This can be achieved in part through increasing parallelization of processing, and a transition of processing hardware from traditional central processing units (CPUs) to graphical processing units (GPUs) and tensor processing units (TPUs). Studies of science application performance at modern supercomputing facilities have demonstrated significant improvements in energy efficiency by using GPUs instead of CPUs for processing, although the impacts vary significantly with the specific application [61]. However, history has repeatedly demonstrated that increased efficiency does not necessarily translates into reduced total energy (Jevons paradox) [62], including in the case of the usage of computing farms.
The dominant concern in the era of de-carbonized power generation will be the global warming impacts associated with the manufacturing process that leads to “embodied carbon” in the computing hardware. This involves the emissions associated with extracting the raw materials (such as heavy metals) that are needed in computers, and the emissions associated with the manufacturing process itself. For instance, the embodied carbon associated with a typical Dell server used in a data center is about 1300 kgCO2[63]. The embodied carbon associated with computer hardware can be reduced through improvements to manufacturing processes, the use of lower-impact materials, and an improved recycling practice of computing components [64]. While these are beyond the control of the end users, particle physicists can seek to reduce the impact of embodied carbon by selecting vendors with the most sustainable practices, by trying to extend the lifetime of existing computing hardware, thereby reducing the rate of replacement, and by adhering to strict recycling practices.
@@ 5 EMISSIONS ASSOCIATED WITH TRAVEL AND OTHERS
We use this information as inputs to a simple model (see Table 4) of the carbon emissions associated with travel to the experiment site by collaborators, using the FCC as an example. We assume a collaboration size of 5000, in which half of the collaborators reside near CERN, the experiment site. One quarter are “regional” participants who work London-like distances from CERN and visit the experiment four times per year. The remaining quarter are “remote” participants who live Chicago-like distances from CERN and visit the experiment only two times per year, because of the large distance. The regional participants produce about 1.7 kg CO2e/year from their air travel, while the remote participants produce about 8.4 kg CO2e/year. In one year, the collaboration produces 12.6 kt CO2e emissions from air travel, which summed over a 20-year physics program is 0.25 Mt CO2e. Comparing to the estimates of Section 2, this amount is more than that associated with the FCC accelerator components, and in fact about one quarter of the emissions associated with the civil construction. This makes travel a subleading but still significant contribution to the carbon emissions over the lifetime of the project.
@@
International conferences also have the potential to be a significant source of carbon emissions, as they draw people from all over the planet, many of whom arrive by long-haul airplane flights. This is issue is not unique to future colliders, and we refer readers to previous discussions of the topic [70].
@@ 6 OBSERVATIONS AND R&D PRIORITIES
Research and development activities focused on minimizing the environmental impact of future colliders have been in progress for a few years. Several presentations in the 2024 editions of the Sustainable HEP conference [71] and of the ICHEP conference [72] reported on those efforts. Of note was the announcement of the start of the iSAS (innovate for Sustainable Accelerating Systems) [73] project funded by EU Horizon 2020. This initiative plans to explore three key innovation directions for energy savings: in the RF system, the cryogenics, and energy savings from the beam itself [74]. In addition, the CEPC collaboration showed interesting results regarding the power savings which could be achieved using the mid-temperature baking technology, which can enhance the quality factor of the RF cavities by a factor of five compared to the current electropolishing technique [31]. Finally, more projects are using the LCA method to optimize their designs to reduce their environmental impact, as demonstrated by the ISIS-II project [75].
The accelerator construction analysis described in Section 2.1 suggests that the emissions associated with the cement needed for the tunnel could easily dominate the total emissions of a future accelerator, and thus is in the greatest need of mitigation. Consequently, any new accelerator design which minimizes the length of the tunnels has a clear advantage from a carbon emissions and potentially a cost perspective. The recent hybrid, asymmetric, linear Higgs factory design based on plasma-wakefield and RF acceleration is such an example [76]. As mentioned by the engineering companies involved in the various LCA estimates available, the choice of material used for the civil construction can result in a wide range of embedded carbon. As an example, in the standard ICE database [28] there are 23 different types of cement, which is responsible for the high embedded carbon of concrete. The standard cement assumed by most LCAs is CEM I, Ordinary Portland Cement, with a carbon intensity of 0.912 kgCO2e/kg. As explained in [30], if this cement is mixed in with some Supplementary Cementitious Materials (SCMs), such as fly ash, Ground Granulated Blast furnace Slag (GGBS), limestone powder, or calcined clay, then the carbon intensity may be reduced. For example, the CEM II/B-S - 28% GGBS which includes between 21-25% of GGBS material has a carbon intensity of 0.672 kgCO2e/kg, a reduction of 26% compared to CEM I. However, one has to consider the potential drawbacks of using those alternative materials. These include whether they would be appropriate for a structure meant to have a lifetime greater than sixty years, and what the cost implications would be, given that the materials are becoming are in increasing demand due to sustainability concerns worldwide. Finally, there are also very important R&D avenues being investigated in sequestering carbon during the cement production stage [77].
@@
As discussed in section 4, emissions from computing originate from the embedded carbon in the computing components and from the electricity used to power those components. For the latter, as mentioned in section 2.3, western countries expect to have a decarbonized electricity grid by 2040, hence a significant fraction of computing farms that will be used to process the energy-frontier collected data will have negligible indirect emissions. However, as already mentioned, demand for electricity is expected to grow significantly due to the necessity for countries to achieve their Net Zero target by 2050. For example, in the U.K., the Climate Change Committee forecasts that electricity demand will double by 2050 compared to 2018 levels [33], with a similar picture in the U.S. [79]. In addition, installed capacity needs to be at least a factor of two larger than the peak demand for electricity. Achieving this target while only using renewable or nuclear sources will be a significant challenge and thus it is reasonable to assume that the price of electricity will either remain the same as today or more likely increase. As a consequence, it is of vital importance for computing farms and for colliders used for particle physics projects to use as little power and total electrical energy as possible. An interesting area of computing R&D studies is frequency modulation, where the CPU clock can run at different rates and be dynamically adjusted to adapt to the power demand of the grid [80].
@@
Some authors have suggested that one possible avenue to ensure that renewable electricity sources are used for the power needs of future facilities is for those facilities to purchase such renewable providers (e.g. solar farms or wind turbine farms). It has also been suggested that CERN could get its electricity from solar farms in North Africa, transmitted via sub-sea cables (see Case Study 3.3 in [8]). First it is important to note that Scope 2 (indirect) emissions do not include the emissions associated with the construction and the embodied carbon of the electricity sources, while if a facility builds or purchases such a source, those emissions would need to be included in their Scope 1 emissions. Finally, we believe that the implications of particle physics research facilities acquiring such resources should be explored and discussed. Given the climate emergency facing humanity, particle physics researchers should instead contribute to the work needed for every country to decarbonize their electricity grid, supporting the minimization of Scope 2 emissions for particle physics projects.
@@
As discussed in section 5, one way to mitigate travel emissions is to maximize the use of videoconferencing systems. However, as experienced during the COVID-19 pandemic, this does not capture the spontaneous interactions crucial to the exchange of ideas, and it can also lead to online fatigue. The field of particle physics has a long tradition of being an innovator in collaborative tools, such as the invention of the World Wide Web. This spirit of innovation should continue to enable minimum travel. This will be especially crucial given that decarbonizing flight is a very difficult goal to achieve and, some countries have Green Parties advocating legislation aimed at introducing a frequent-flyer levy [81]. The field of particle physics could champion a vast expansion of the use of Virtual/Augmented Reality headsets. These can help make online social interactions more natural and also support participation from the Global South in particle physics projects.